OpenStudy (anonymous):

how to integrate x^2/(1+x^3)

OpenStudy (anonymous):

i will show you a trick to solve this problem

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

as you can see the numerator holds the derivative of the denomirator

OpenStudy (anonymous):

the answer would be

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\ln(1+x^{3})\]

OpenStudy (anonymous):

how did this happen?

OpenStudy (anonymous):

The problem is what we're the steps?

OpenStudy (anonymous):

yes i will show you the steps

OpenStudy (anonymous):

but i want you to visualize what is going on

OpenStudy (anonymous):

before i show u the steps

OpenStudy (anonymous):

Thanks I've been trying or two hours

OpenStudy (anonymous):

please bear in mind a fact in calculus when you see the derivative at the numerator of the denominator then it is a natural log

OpenStudy (anonymous):

please let me know if you don't understand what i mean

OpenStudy (anonymous):

step one: let u = \[(1+x^{3})\]

OpenStudy (anonymous):

Yes I understand

OpenStudy (anonymous):

step two: d.w.r.x -> \[\frac{ du }{ dx }=3x ^{2}\]

OpenStudy (anonymous):

this means that : \[dx = \frac{ du }{ 3x ^{2} }\]

OpenStudy (anonymous):

Ok so ffar

OpenStudy (anonymous):

step three: look at the main function you are trying to integrate which is : \[\int\limits_{}\frac{ x^2dx }{ 1+x^3 }\]

OpenStudy (anonymous):

now put that in terms of the u

OpenStudy (anonymous):

\[\int\limits\limits_{}\frac{ x^2du }{ 3x^2u }\]

OpenStudy (anonymous):

now tell me do you see what is going on?

OpenStudy (anonymous):

\[\frac{1}{3}\int\limits\limits\limits_{}\frac{ du }{ u }\]

OpenStudy (anonymous):

Yes so far

OpenStudy (anonymous):

\[\frac{1}{3}\ln (u) + c\]

OpenStudy (anonymous):

now try yourself to put this last equation in terms of x

OpenStudy (anonymous):

Yes I see it :)

OpenStudy (anonymous):

recall \[\frac{1}{3}\ln (1+x^3) + c\]

OpenStudy (anonymous):

Sorry couldn't type it in right

OpenStudy (anonymous):

so now you should be able to work out by head

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

i gave the answer

OpenStudy (anonymous):

Yes I was so confused on how the 1/3 was coming in front of the ln

OpenStudy (anonymous):

I can't thank you enough

OpenStudy (anonymous):

its ok if you have another question i can answer it hopefully

OpenStudy (anonymous):

you can give me a medal hehehe

OpenStudy (anonymous):

For now that's it , but are you usually online? I might have a question later

OpenStudy (anonymous):

not really not always

OpenStudy (anonymous):

Of course you deserve a hundred medals..just how do I give u one

OpenStudy (anonymous):

i only was here since yesterday but there are others who can help as well

OpenStudy (anonymous):

i don't know hehehe

OpenStudy (anonymous):

Il find out..ur getting the medal

OpenStudy (anonymous):

thanx that is kind of you

OpenStudy (anonymous):

you are a clever girl you pick up quick the answer

OpenStudy (anonymous):

It was part of my Differential equation hw..I was stuck on the integration

OpenStudy (anonymous):

what year are you?

OpenStudy (anonymous):

Junior but haven't taken calc in over 5 years!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

man too many questions on this forum

OpenStudy (anonymous):

right

OpenStudy (anonymous):

You? Math major

OpenStudy (anonymous):

no i am a communication engineer

OpenStudy (anonymous):

Cool well best of luck and thanks again :)

OpenStudy (anonymous):

you too all the best i will see who can i help here